A sports car accelerates from rest to 90km/h in 4.1s. What is the average acceleration in m/s^2?
Answer #1

Average acceleration = (Change in velocity) /(Time interval)

aavg = (v2-v1) / (t2 - t1)

initial velocity v1 = 0 m/s

final velocity  v2   = 90 km/hr

conversion 1 km/hr = 5/18 m/s

therefore final velocity  v2   = 90 km/hr = 90 *(5/18 ) = 25 m/s

time interval (t2 - t1) = 4.1 seconds

therefore aavg = (v2-v1) / (t2 - t1) = (25-0)/4.1 = 6.09756 m/s2 = 6.1 m/s

Answer is 6.1 m/s2